# AI Gender Bias, Disparities, and Fairness: Does Training Data Matter?

### 17 Dec 2023 v1

@@ 1. Introduction
Artificial Intelligence (AI) has permeated various aspects of human life, revolutionizing industries and practices with its advanced computational capabilities. However, the potential of AI is often marred by inherent biases, especially gender biases, that can have profound implications on the fairness and effectiveness of AI applications. Gender bias in AI, particularly in the context of automatic scoring systems, poses a significant challenge, echoing societal stereotypes and disparities (Zhai and Nehm, 2023). Studies such as (Leavy, 2018; Nadeem et al., 2020; Cirillo et al., 2020) have highlighted the critical need for diversity and gender theory in machine learning to address these biases.
@@
Pseudo-AI bias, a concept introduced by Zhai and Krajcik (Zhai and Krajcik, 2022), counters the mechanical bias inherent in AI systems and often draws unnecessary anxiety about AI. The mechanical bias, while not intentional, can lead to discriminatory outcomes, especially in gendered contexts. The work of Bolukbasi et al. (Bolukbasi et al., 2016) and Hall et al. (Hall and Ellis, 2023) further underscores the impact of such biases stemming from word embeddings and supervised machine learning models, respectively.
@@
The significance of this study lies in its focus on providing statistical evidence regarding the gender-neutral nature of the mixed-gender trained model by analyzing more than 1000 human-graded student responses across different gender categories. We have trained models using male-only, female-only, and mixed written datasets and evaluated the performance of these models against males and females only and mixed testing data. This research employs standard techniques to measure gender bias using scoring accuracy difference (pair t-test (Mowery, 2011)), gender disparity using mean score gaps between predicted scores for different genders (Wang and Brown, 2007; Wilson et al., 2023), and fairness using Equalized Odds (Hardt et al., 2016), to evaluate gender bias, discrepancies, and fairness (See Fig 1 for a detailed overview of the approach). The findings reveal a promising trajectory toward analyzing gender issues, evidenced by the insignificant accuracy difference that means non-biased prediction, low MSGs mean non-disparate prediction, and minimal odds gradient gamma (¡0.01) mean fair prediction for mixed and separate models.
This study contributes to the ongoing discourse on gender bias in AI, aligning with recent research (Franzoni, 2023; Lima et al., 2023; Hall and Ellis, 2023) that emphasizes the need for gender-sensitive AI development. Furthermore, it resonates with the efforts in mitigating gender bias through comprehensive reviews and algorithmic adjustments (Sun et al., 2019; Li et al., 2022; Lu et al., 2020).
The motivation behind this study is rooted in the ethical imperative to ensure fairness and equity in AI systems, a sentiment echoed by Slimi (Slimi, 2023) and O’Connor & Liu (O’Connor and Liu, 2023). By demonstrating that targeted mix model training doesn’t have traces of gender bias, this research offers a blueprint for developing more equitable AI systems in education and beyond, contributing to the broader discourse on ethical AI (Nemani et al., 2023; Manresa-Yee and Ramis, 2021).
@@ 2.1. AI in Education
Integrating AI in education marks a paradigm shift in teaching and learning. AI’s role in education encompasses a broad spectrum, from personalized learning to administrative efficiency and pedagogical innovations.
Roles and Impacts: AI’s implementation in educational contexts is multifaceted. Studies (Holmes et al., 2021; Berendt et al., 2020) highlight the ethical dimensions of AI in education, emphasizing the need for frameworks that balance technological advancement with moral and social values. Felix (Felix, 2020) and Qadir (Qadir, 2023) discuss the evolving role of AI, from augmenting teaching methodologies to supporting administrative tasks, offering insights into how AI can enhance the educational experience for both teachers and students.
Educational Analytics and Assessment: AI-driven analytics play a crucial role in assessing student performance and identifying learning gaps. By processing large volumes of data, AI systems can provide educators insights into student progress, enabling more informed decision-making and targeted instructional strategies.
Ethical and Social Considerations: The incorporation of AI in education also raises important ethical questions. Data privacy, algorithmic fairness, and the digital divide are central to the discourse on AI in education. A recent study (Holmes et al., 2021) underscores the importance of addressing these ethical concerns, advocating for a community-wide framework that ensures AI is used responsibly and equitably in educational settings.
AI’s role in education rapidly evolves, offering exciting possibilities for enhancing teaching and learning experiences. However, the successful integration of AI in education hinges on addressing the ethical, social, and practical challenges accompanying these technological advancements.
@@ 2.2. Automatic Scoring in Education
The advent of AI in education has brought significant advancements in automatic scoring systems, revolutionizing how student assessments are conducted and analyzed (Zhai et al., 2020b, a; Latif and Zhai, 2023a). This technology offers scalability and efficiency for instruction and introduces new paradigms in educational assessment practices.
Current State and Advancements: González-Calatayud et al. (González-Calatayud et al., 2021) provide a comprehensive systematic review of AI applications in student assessment, highlighting various methodologies and their effectiveness. Zhai et al. (Zhai et al., 2021a) reviewed AI-based assessments in science and discuss the state of the art and practice, focusing on the latest advancements and research findings. Their work illustrates how AI technologies, particularly machine learning and natural language processing, have effectively evaluated student responses in diverse subjects, though with varying degrees of scoring accuracies.
Challenges in Implementation: Despite these advancements, implementing automatic scoring is challenging. Studies (Zhai et al., 2022) delve into the application of machine learning in scientific assessments, shedding light on the complexities and nuances involved in accurately assessing student models. Zhai & Nehm (Zhai and Nehm, 2023) further explore these challenges, particularly in formative assessments, and discuss how AI can be optimized for more effective educational outcomes.
Machine-Human Score Agreements: A critical aspect of automatic scoring is the agreement between machine-generated scores and human evaluations. A meta-analysis of machine learning-based science assessments was conducted to understand the factors impacting machine-human score agreements (Zhai et al., 2021b). Their findings suggest that while AI can closely mimic human scoring in certain contexts, discrepancies still exist, underscoring the need for continuous refinement of AI algorithms in educational settings.
Potential and Future Directions: The potential of AI in automatic scoring extends beyond efficiency and scalability. It offers opportunities for personalized feedback, adaptive learning environments, and a deeper analysis of student understanding (Zhai, 2021; Lee et al., 2023b). As suggested by researchers, the future of AI in education lies in generative AI’s ability to complement and enhance traditional teaching methods, providing valuable insights into student learning processes and outcomes. LLMs represent a transformative tool with immense potential. Researchers have leveraged these models in automatic scoring of science writing (Latif and Zhai, 2023b; Lee et al., 2023a).
@@ 2.3. AI Gender Bias, Disparities, and Fairness
Gender bias, disparities, and fairness in AI, particularly in education, is a topic shrouded in both myths and harsh realities. We consider the systematic definition of bias for AI (similar to (Larrazabal et al., 2020)), which refers to systematic rather than random error in which we analyze the calculated error in scoring accuracy of trained AI models. Gaps in automatically predicted machine scores for responses written by different genders of students are considered gender disparities in AI (concept defined by OECD (for Economic Co-operation and, OECD)). Further, AI fairness as equity (Holstein and Doroudi, 2019) is considered where the AI model assumes to be fair if it demonstrates an equal proportion of false positives and true positives for automatically predicted scores of different gender written responses. This section aims to dissect these aspects, providing a clearer understanding of the extent and implications of gender issues in AI systems.
Unraveling Myths: Common myths around AI gender bias often stem from misunderstandings of AI functionalities and oversimplified interpretations of gender disparities. One such myth is the belief that AI systems are inherently unbiased as they are based on algorithms and data. However, as Bolukbasi et al. (Bolukbasi et al., 2016) and Hardt et al. (Hardt et al., 2016) illustrate, biases in training data and algorithmic design can lead to gender-biased outcomes. Another myth is that gender bias in AI is solely a technical issue that can be resolved with more data or better algorithms. While these are crucial factors, as pointed out by Lu et al. (Lu et al., 2020) and Manresa-Yee & Ramis (Manresa-Yee and Ramis, 2021), addressing gender bias in AI also requires a broader understanding of societal and cultural contexts.
Confronting Realities: The reality of gender bias in AI is multifaceted and deeply rooted in both the technical and social domains. The concept of pseudo-AI bias (Zhai and Krajcik, 2022) highlights how biases can be subtly embedded in AI systems, often reflecting societal stereotypes. These biases can manifest in various forms, such as skewed representations in datasets or biased interpretations of AI outcomes.
Impact on Education: In educational settings, AI gender bias can have significant consequences. It can lead to unfair assessment practices, reinforce gender stereotypes, and create barriers to equal educational opportunities. Madaio et al. (Madaio et al., 2022) delve into the structural injustices that AI can perpetuate in educational contexts, advocating for a more holistic approach to understanding and addressing these biases.
Debunking Myths and Embracing Realities: Ultimately, debunking myths and embracing the realities of AI gender bias in education is crucial for developing AI systems that are fair, equitable, and beneficial for all learners. This calls for a multi-disciplinary approach incorporating insights from gender studies, ethical AI practices, and educational theory to ensure that AI is a tool for empowerment rather than perpetuating existing disparities.
Understanding and addressing AI gender bias in education requires moving beyond myths and confronting the complex realities of how gender biases are embedded and perpetuated in AI systems. This understanding is crucial for developing AI applications that are equitable and effective in diverse educational contexts.
@@
Study’s Approach to Bias, Disparities, and Fairness Analysis: The study employs a rigorous methodology to analyze gender issues in automatic scoring. Utilizing a fine-tuned version of BERT (similar to (Liu et al., 2023)) and GPT-3.5 (similar to (Latif and Zhai, 2023b)), the research analyzes around 1000 human-graded student responses, proportionally distributed among male, female, and mixed datasets across six assessment items. The study measured bias by pair t-test of scoring accuracy between different gender-specifically trained models (Mowery, 2011); the gender disparities by mean score gaps by gender(Wang and Brown, 2007; Wilson et al., 2023); and fairness by EO (Hardt et al., 2016). These methods comprehensively show how gender issues manifest in AI scoring systems for mixed and separately trained models.
@@
Gender-Neutral Nature of Mixed Trained Models: The core hypothesis of the study is that training AI models with mixed-gender datasets doesn’t show gender preferences for automatic scoring upon testing male and female responses. This hypothesis is grounded in the understanding that diverse training data can help AI algorithms learn more nuanced, gender-neutral patterns, thus mitigating biases that arise from unbalanced or skewed datasets.
Setting Ground to Avoid AI gender issues: A critical aspect of this study is its commitment to dispelling myths surrounding AI and gender issues. By employing a data-driven, empirical approach, the study demonstrates that gender issues in AI are not immutable and can be addressed through thoughtful and deliberate training methods. This counters the common misconception that AI systems are inherently biased or that biases in AI are too complex to be resolved.
Implications for AI Development: The findings of this study have significant implications for AI development, particularly in educational contexts. They suggest carefully considering training data and applying targeted analysis techniques can lead to more equitable and fair AI systems. This approach aligns with the broader discourse on ethical AI development, emphasizing the need to incorporate diversity and gender theory in AI training practices.
To sum up, this study contributes to the ongoing efforts to analyze many gender issues in AI through targeted training and rigorous analysis. By providing empirical evidence and a methodological framework, it sets the ground for developing AI systems that are both technically sound and ethically responsible, thereby avoiding common myths.
@@ 3. Methodology
This study aims to investigate gender issues in AI-based scoring systems based on State-Of-The-Art Language Models BERT and GPT-3.5 models. Large Language models tend to revolutionize education technology (Latif et al., 2023) and raise potential challenges such as their bias, disparity, and fairness, which need to be addressed. A comprehensive methodology was adopted, comprising different training and testing phases for models and employing various statistical techniques to evaluate gender bias, disparities, and fairness.
@@ 3.1. Dataset
The study utilized a meticulously categorized dataset of student-written responses to six assessment items, each classified under the multi-class category. The datasets were categorized based on reported gender (i.e., male and female; we eliminated the data that did not report gender information), and a mixed dataset incorporated responses from both genders. Based on the availability of data, we made use of the maximum of available data while maintaining the ratio between training and testing data for each item by gender. Consequently, we randomly identified three training datasets for each item: mixed, male, and female training datasets; and three testing datasets for each item. The detailed composition of each assessment item’s dataset is presented in Table 1.
Dataset Overview:
  • Falling Weights: Includes 1148 training and 230 testing samples in the mixed dataset, with 342 training and 87 testing samples for Males and 478 training and 120 testing samples for females.
  • Gelatin: Comprises 918 training and 230 testing samples in the mixed dataset, with 344 training and 86 testing samples for Males and 477 training and 120 testing samples for females.
  • Bathtub: Contains 916 training and 230 testing samples in the mixed dataset, with 364 training and 92 testing samples for Males and 450 training and 113 testing samples for females.
  • Sandwater1 and Sandwater2: Both have similar structures, with Sandwater1 having 915 training and 229 testing samples and Sandwater2 having 913 training and 229 testing samples in the mixed dataset. For both, Males contributed 364 training and 92 testing samples, while Females contributed around 450 training and 113 testing samples.
  • Two Boiling Situations: Consists of 909 training and 228 testing samples in the mixed dataset, with 363 training and 91 testing samples for Males and 448 training and 113 testing samples for females.
@@ 3.4. Mean Score Gap
Mean Score Gap (MSG) was utilized to evaluate the disparity in scores between different genders. MSG can be generated either by human experts or machine algorithmic models. Given that human-assigned scores have undergone rigorous scrutiny, we used MSG generated by human-assigned scores as the standard to evaluate machine-generated MSG. It can be seen that if machine-generated MSG is larger than human-generated MSG for the same testing data, the machine exacerbates the gender difference and may potentially create inequity. Wang and Brown (Wang and Brown, 2007) and Wilson et al. (Wilson et al., 2023)highlighted the importance of such comparisons in their study on automated essay scoring and automatic scoring of written scientific arguments. MSG is calculated as:

### 26 Dec 2023 v2

@@ 1. Introduction
Today’s AI is so advanced in computation that it has penetrated deep into the well of human life, shaking up industries and practices. But the possibility of artificial intelligence is always spoiled by built-in biases, especially gender biases, which often cause serious problems. In particular, discrimination against women in AI takes the form of influential automatic scoring systems that reproduce societal stereotypes and gaps (Zhai and Nehm, 2023). Among others, see (Cirillo et al., 2020; Leavy, 2018; Nadeem et al., 2020) have stressed the important role that diversity is certain to play in machine learning over these questions of bias.
@@
Zhai and Krajcik (2022) introduce the concept of pseudo-AI bias to combat this mechanical AI bias, often giving rise to unnecessary fear about artificial intelligence. Although unintentional, the mechanical bias leads to discriminatory outcomes. These become especially glaring where gender is involved. Such biases in exceedingly widely used word embeddings and supervised machine learning models are further underscored by the work of Bolukbasi et al. (2016) and Hall and Ellis (2023).
@@
This is the significance of this study. By analyzing over 100 human-graded questions across various gender categories, it uses statistical evidence to show that both mixed-trained models are gender-neutral. Based on our male-only, female-only, and mixed written datasets (training), we have tested the performance of these models, creating both males and females or combined testing data. We use standard methods such as scoring accuracy difference (pair t-test (Mowery, 2011)) to measure potential gender bias, mean score gaps (Wang and Brown, 2007; Wilson et al., 2023)between machine and human scores for different genders to measure gender disparity, and Equalized Odds (Hardt et al., 2016) to measure fairness. (See Figure 1 for a detailed overview of the approach). The findings reveal a promising trajectory toward analyzing gender issues, evidenced by the insignificant accuracy difference that means non-biased prediction, low MSGs mean non-disparate prediction, and minimal odds gradient gamma (<.$<.01$) mean fair prediction for mixed and separate models.
Thus, this study adds to the current debate over gender bias in AI. It chimes with recent studies (Franzoni, 2023; Hall and Ellis, 2023; Lima et al., 2023) insisting on developing gender-aware artificial intelligence. Moreover, it is in tune with attempts to remove gender bias through all-round reviews and refinements of algorithms (Li et al., 2022; Lu et al., 2020; Sun et al., 2019).
According to Slimi (2023) and O’Connor and Liu (2023), this study is due to an ethical mandate of fairness or even equity in AI systems. This research shows that targeted mix model training lacks traces of gender bias, providing a pathway for creating more equitable AI in education and other contexts. It contributes to the growing debate about ethical or socially beneficial artificial intelligence AI (Manresa-Yee and Ramis, 2021; Nemani et al., 2023).
@@ 2.1. AI in Education
AI, as an integral part of education, represents a fundamental change in the teaching and learning process. From personalized learning to administrative efficiency and pedagogical innovations, the fields in which AI can play a role are very diverse.
Roles and Impacts: Multifaceted implementation of AI in educational contexts. There are also studies (Holmes et al., 2021; Berendt et al., 2020) that discuss the ethical risks of AI in education and insist on keeping pace with technological progress while not violating moral or social values. How does AI fit into the educational process? Felix (2020) and Qadir Qadir (2023) share their perspectives on this, with changing applications of artificial intelligence from fortifying teaching methodologies to assisting administrative work.
Educational Analytics and Assessment: AI-based analysis is vital for evaluating student performance and discovering learning deficiencies. AI systems capable of handling large amounts of data can help educators interpret how students are doing so that they make more informed decisions and plan better-targeted approaches.
Ethical and Social Considerations: AI in education also raises important ethical questions. Algorithmic fairness, data privacy, and the digital divide are key to discussing AI in education. One recent study (Holmes et al., 2021) shares these ethical concerns, urging a community-wide framework to ensure that AI is used responsibly and equitably in educational settings.
The educational use of AI is developing rapidly, bringing new promise for enriching pedagogy and learning. Yet achieving the integration of AI in education depends on meeting the ethical, social and practical challenges posed by these technological developments.
@@ 2.2. Automatic Scoring in Education
Now that the era of AI in education is arriving, automatic scoring systems have evolved by leaps and bounds (Zhai et al., 2020b, a; Latif and Zhai, 2023a). Today’s student assessments differ greatly from those of yesteryear. It provides scalability and efficiency for training while promoting new paradigms in educational tests.
Current State and Advancements: In a recent and detailed systematic review of studies on AI applications in creating student assessments, González-Calatayud et al. (2021) cover methodologies that have been used with information about their effectiveness from different countries around the world in October 2015, Zhai et al. (2021a) had published a review of AI-based assessments on science that covered the state-of-the-art and practice concerning their latest developments as well as research outcomes. Their work shows that various AI technologies, especially machine learning and natural language processing, have already been proven to evaluate student answers in different subjects accurately.
Challenges in Implementation: These developments notwithstanding, making automatic scoring a reality is difficult. Zhai et al. (2022) explore the use of machine learning in scientific assessments, providing insight into how difficult it is to evaluate student models accurately. Zhai and Nehm (2023) look further into these problems, especially as they pertain to classroom assessments of student performance. They also consider how AI can be designed better for improved educational outcomes.
Machine-Human Score Agreements: Agreeing scores between machines and men is an important factor in automatic scoring. Zhai et al. (2021b) conducted a meta-analysis of machine learning-based science assessments to learn the reasons for low agreements between machine and human scores. Their conclusions indicate that although, in some aspects, AI can come extremely close to the humanly scored version, differences nonetheless remain. There is still a need to review and improve individual algorithm parameters within educational environments continually.
Potential and Future Directions: Improving efficiency and scalability is not all AI can do for automatic scoring. It provides room for personal feedback, an adapting learning environment, and an in-depth assessment of student comprehension (Lee et al., 2023b; Zhai, 2021). According to researchers, the future of AI in education is for generative AI technology to complement and augment verbal teaching methods. It could go further than that by even allowing us glimpses into student learning processes and outcomes like never before possible. With great potential, LLMs are a revolutionary tool. These models have been used by researchers in grading science writing automatically (Latif and Zhai, 2023b; Lee et al., 2023a).
@@ 2.3. AI Gender Bias, Disparities, and Fairness
Gender bias, disparities, and fairness in AI–especially educationally speaking, is a subject clouded by half-truths and hard facts. We consider the systematic definition of bias for AI (similar to (Larrazabal et al., 2020)), which refers to systematic rather than random error in which we analyze the calculated error in scoring accuracy of trained AI models. Gaps in automatically predicted machine scores for responses written by different genders of students are considered gender disparities in AI (concept defined by OECD (for Economic Co-operation and, OECD)). Further, AI fairness as equity (Holstein and Doroudi, 2019) is considered where the AI model assumes to be fair if it demonstrates an equal proportion of false positives and true positives for automatically predicted scores of different gender written responses. This section aims to dissect these aspects, providing a clearer understanding of the extent and implications of gender issues in AI systems.
Unraveling Myths: Misunderstanding Missing Information and flawed conclusions are almost always behind common myths about the gender bias of AI. One such myth is the belief that AI systems are naturally unbiased since they rely on algorithms and data. But as Bolukbasi et al. (2016) and Hardt et al. (2016) show, biases are inherent not only to the training data on which algorithmic design is based but even to its mechanisms of operation; they can result from seemingly gender-neutral words or expressions having a greater affinity with the Another myth is that the gender bias in AI lies wholly at a technical level, where it depends on solving these difficulties by gathering more data or writing better algorithms. These are essential factors, as recognized by Lu et al. (2020) and Manresa-Yee and Ramis (2021). However, overcoming gender bias in AI requires a more profound awareness of societal structures and cultural values.
Confronting Realities: Gender bias in AI is complex and multilayer, both quantitatively technical and qualitatively social. A type of bias is pseudo-AI Bias Zhai and Krajcik (2022) that reflects society’s stereotypes. Bias can appear in various forms, such as distorted dataset representations or biased interpretations of results from AI.
Impact on Education: In educational settings, this bias affects outcomes. It means unequal assessment, reinforced gender stereotypes, and barriers to genuine educational opportunity. Madaio et al. (2022) analyze structural injustices that may be propagated through AI applications in educational contexts. They call for more comprehensive approaches to understanding these biases.
Debunking Myths and Embracing Realities: The bottom line is that both in education and elsewhere, slaying myths about AI gender bias – along with being aware of it– can lead to developing fairer, more equitable systems for all learners. This requires an integrated, multi-faceted approach that incorporates the ideas of gender studies experts and people from practice in ethical AI design and educational theory to ensure our applications are not simply making existing dignities unequal. To understand and resolve AI gender discrimination in education, we must abandon myths and face up to the actual, much more complicated ways that various biases both exist within and arise from systems of artificial intelligence. Such an understanding is critical to creating AI applications that are fair and effective in different educational settings.
@@
Study’s Approach to Bias, Disparities, and Fairness Analysis: The study employs a rigorous methodology to analyze gender issues in automatic scoring. We have fine-tuned standard BERT (similar to (Liu et al., 2023)) and GPT-3.5 (identical to (Latif and Zhai, 2023b)), using more than 1000 human expert’s rated student responses, proportionally distributed among male, female, and mixed datasets across six assessment items. The study measured bias by pair t-test of scoring accuracy between different gender-specifically trained models (Mowery, 2011); the gender disparities by mean score gaps by gender (Wang and Brown, 2007; Wilson et al., 2023); and fairness by EO (Hardt et al., 2016). These methods comprehensively show how gender issues manifest in AI scoring systems for mixed and separately trained models.
@@
Gender-Neutral Nature of Mixed Trained Models: The study’s core hypothesis is that AI models trained using mixed-gender datasets don’t demonstrate gender prejudice in automatic scoring when tested on male and female responses. At its core is the notion that AI algorithms can only learn more gender-neutral patterns with a greater diversity of training data, thus eliminating biases resulting from unbalanced or skewed datasets.
Setting Ground to Avoid AI gender issues: A critical element of this study is its aim to demystify AI and gender problems. The study uses a data-driven, empirical approach to show that difficulties in AI surrounding gender are not immovable. Through careful and reasoned training methods, they can be solved. This goes against the prevailing misconception that all AI systems are biased or that arriving at bias-free AI is too complicated.
Implications for AI Development: These findings are significant for AI development in educational environments. They say that training data must be carefully thought about, and applying an appropriate analysis technique can result in fairer, more equitable AI systems. This is consistent with the broader debate on mainstreaming ethics in AI development, which stresses that all aspects of theory should be used to train AIs.
In short, this study adds to the ongoing efforts of many gender issues in AI through targeted training and rigorous analysis. It also provides empirical evidence and a conceptual framework, paving the way for adequate technical knowledge and full moral responsibility to develop responsible AI systems that do not fall prey to many wrong beliefs.
@@ 3. Methodology
This study aims to investigate gender issues in AI-based scoring systems based on state-of-the-art Language Models BERT and GPT-3.5 models. Large Language models tend to revolutionize education technology (Latif et al., 2023) and raise potential challenges, such as bias, disparity, and fairness, which must be addressed. A comprehensive methodology was adopted, comprising different training and testing phases for models and employing various statistical techniques to evaluate gender bias, disparities, and fairness. Figure 1 presents an overview of the research design.
@@ 3.1. Dataset
The study utilized a meticulously categorized dataset of student-written responses to six assessment items, each classified under the multi-class category. The datasets were categorized based on reported gender (i.e., male and female; we eliminated the data that did not report gender information), and a mixed dataset incorporated responses from both genders. Based on the availability of data, we used the maximum of available data while maintaining the ratio between training and testing data for each item by gender. Consequently, we randomly identified three training datasets for each item: mixed, male, and female training datasets and three testing datasets for each item. The detailed composition of each assessment item’s dataset is presented in Table 1.
@@ 3.4. Mean Score Gap
Mean Score Gap (MSG) was utilized to evaluate the disparity in scores between different genders. MSG can be generated either by human experts or machine algorithmic models. Given that human-assigned scores have undergone rigorous scrutiny, we used MSG generated by human-assigned scores as the standard to evaluate machine-generated MSG. It can be seen that if machine-generated MSG is larger than human-generated MSG for the same testing data, the machine exacerbates the gender difference and may potentially create inequity. Wang and Brown (2007) and Wilson et al. (2023)highlighted the importance of such comparisons in their study on automated essay scoring and automatic scoring of written scientific arguments. MSG is calculated as: